OpenStudy (anonymous):

Explain how you would graph -0.5x + 0.1y = 1.2.

OpenStudy (anonymous):

Solve for y, and then that will give you the slope and y intercept.

OpenStudy (anonymous):

But, I would first multiply it by 10 to get rid of the decimals.

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

Multiply (-0.5x + 0.1y = 1.2)*10. You get -5x+y=12

OpenStudy (anonymous):

now solve -5x+y=12 for y. so y=5x+12. The slope is 5x and y intercept is 12.

OpenStudy (anonymous):

ohh ok thank youu

OpenStudy (anonymous):